OpenStudy (anonymous):

Solve for c. ab + c = 5

TheSmartOne (thesmartone):

First subtract ab on both sides what do you get?

OpenStudy (gabylovesyou):

ab + c = 5 ab - ab + c = 5 - ab c = 5 - ab @TheSmartOne ?

TheSmartOne (thesmartone):

Correct @Gabylovesyou

OpenStudy (gabylovesyou):

yay me ^-^ @candy_cano111 did u understand how i did it ?

OpenStudy (anonymous):

@Gabylovesyou yes thank you! (: